# Prompt Stellar and Binary Black Hole Mergers from Chemically Homogeneous Evolution in Triples

### v1

@@
22institutetext: School of Physics and Astronomy, Monash University, Clayton, VIC 3800, Australia 33institutetext: OzGrav: Australian Research Council Centre of Excellence for Gravitational Wave Discovery, Clayton, VIC 3800, Australia 44institutetext: Institute for Astronomy, School of Physics, Zhejiang University, 310058 Hangzhou, China 55institutetext: Center for Computational Astrophysics, Flatiron Institute, New York, NY 10010, USA 66institutetext: National Astronomical Observatory of Japan, National Institutes of Natural Sciences, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan
Prompt Stellar and Binary Black Hole Mergers from Chemically Homogeneous Evolution in Triples
A. Vigna-Gómez 11    E. Grishin 2233    J. Stegmann 11    A. Olejak 11    S. A. Popa 11    B. Liu 44    A. S. Rajamuthukumar 11    L. A. C. van Son 55    A. Bobrick 2233    A. Dorozsmai 66
@@
Short-period massive binary stars are predicted to undergo chemically homogeneous evolution (CHE), making them prime candidates for producing binary black holes (BBHs) that may merge within the age of the Universe. Most of these binaries have a tertiary companion, and here we explore how a nearby third body possibly influences this evolutionary channel. Our analysis combines analytic treatments of triple dynamics with insights from detailed stellar evolution models, focusing on the role of the von Zeipel-Lidov-Kozai mechanism while also accounting for tidal and general relativistic apsidal precession. We examine the dynamics of triples at three critical evolutionary stages: the zero-age main sequence, shortly after the main sequence, and at the time of BBH formation. We find that, for triples having outer orbital periods less than 70 d(120 d), the inner binary can merge during(or after) the main-sequence stage, leading to a hydrogen-rich(helium-rich) stellar merger. If a stellar merger is avoided, the inner binary may eventually form a BBH. In mildly hierarchical triples, with outer periods of around 100 d, the tertiary component can trigger a rapid merger of the BBH on timescales comparable to the outer orbital period. Stellar tides play a crucial role in determining the fate of the inner binary in such tight triple systems, as they can suppress the perturbative effects of the third star. When tidal forces damp the oscillations induced by the tertiary, the BBH merger may occur soon after stellar collapse. Notably, these outcomes are not restricted to CHE binaries but can also apply to any BBH formed from stars in tight orbits. Mergers in these systems are characterized by the proximity of a tertiary companion and the presence of recently ejected gas, making them promising candidates for electromagnetic counterparts and gravitational-wave signals influenced by nearby tertiary objects.
@@ 1 Introduction
Stars that are several times more massive than the Sun typically form and evolve in multiple systems (e.g., Offner et al., 2023; Li et al., 2024), where they are orbited by one or more distant stellar companions. In multiple-star systems, the evolution of stars can be influenced by the interplay of mass transfer episodes, explosions, and gravitational dynamics (e.g., Eggleton, 2006; Tauris & van den Heuvel, 2023). Recent discoveries reveal that some systems initially classified as binaries are in fact part of a triple in which the most massive star is the outermost tertiary (e.g., Eisner et al., 2022; Kostov et al., 2024). In triples, the tertiary star can dynamically influence the evolution of the system. In this paper, we examine the role of a tertiary star in “tight” high-mass stellar triples—systems where the tertiary’s outer orbital period is less than 1000 days—with a focus on binaries that experience chemically homogeneously evolution (CHE).
@@
CHE has been proposed in the context of single, rapidly rotating stars, where rotationally induced mixing leads to contraction rather than the more typical expansion observed in non-rotating stars during the main sequence (see Maeder, 1987, and reference therein). CHE is predicted to occur for (initially) massive stars ($\gtrsim 40\ M_{\odot}$) in short orbital period binaries ($\lesssim 3\,\rm d$) at low metallicities ($Z\ll Z_{\odot}$), where tidal forces and minimal mass loss maintain the stars rotating rapidly (e.g., de Mink et al., 2009; Song et al., 2013; Hastings et al., 2020). CHE stars are the predicted progenitors of long gamma-ray bursts (Yoon et al., 2006; Aguilera-Dena et al., 2018). CHE binaries that remain massive and compact throughout their evolution are predicted to form binary black holes (BBHs) that can merge within the age of the Universe (Marchant et al., 2016; Mandel & de Mink, 2016; de Mink & Mandel, 2016; du Buisson et al., 2020; Riley et al., 2021), making them potentially detectable through gravitational waves. Although numerous candidates have been proposed in the literature (see Sharpe et al., 2024, and references therein), there is currently no confirmed detection of a CHE binary from observations of stellar binaries nor gravitational waves from BBH mergers.
@@
Given that $\sim 73\pm 16\%$ of massive ($\gtrsim 16\ M_{\odot}$) main-sequence stars are found in triples and/or quadruples (Moe & Di Stefano, 2017), it is natural to consider CHE binary evolution with a tertiary companion. Tight hierarchical triples with CHE inner binaries were first studied semi-analytically in the context of gravitational-wave sources, where a coplanar triple may undergo two sequential BBH mergers that result in a peculiar, massive black-hole remnant (Vigna-Gómez et al., 2021). CHE was subsequently integrated into rapid population synthesis (Riley et al., 2021), enabling the exploration of its role within a stellar triple system (Dorozsmai et al., 2024). Rapid population synthesis of triples with a CHE binary has shed light on two particular outcomes. First, mass transfer of the tertiary onto the inner binary is common, potentially affecting the evolution of the inner binary (Dorozsmai et al., 2024; Kummer et al., 2025). Second, triple dynamics can dynamically assist a merger in the inner binary (Dorozsmai et al., 2024). A dynamically assisted merger can happen swiftly (e.g., Vigna-Gómez et al., 2022), reducing the system from a triple to a binary (e.g., Preece et al., 2024). This change impacts the number of tight inner massive binaries that would undergo CHE if they evolved in isolation.
@@
In this paper, we present an analysis of the impact of a tertiary on dynamically assisted mergers in a CHE binary. Specifically, we assess the role of the von Zeipel-Lidov-Kozai (ZLK) mechanism (von Zeipel, 1910; Lidov, 1962; Kozai, 1962) in tight triples. In the ZLK mechanism, coherent torques exchange the angular momentum between the inner and outer orbits of a hierarchical triple (Naoz, 2016). The early studies of the ZLK mechanism were done in the test-particle limit, where one member of the inner binary is nearly massless, so that the outer angular momentum significantly exceeds the inner one. In this limit, the mutual inclination required to induce a large inner eccentricity is close to 90 deg (e.g., Naoz et al., 2013; Tremaine, 2023, and references therein). However, as the ratio of the inner to the outer orbital angular momentum increases, the critical inclination varies (Anderson et al., 2017; Hamers, 2021). Systems with mild hierarchies, where the ratio of the inner and outer orbital period is not too small, can also achieve larger eccentricities (Antonini & Perets, 2012; Antognini, 2015; Luo et al., 2016; Grishin et al., 2017, 2018b). Nonetheless, additional sources of apsidal precession, such as short range forces, may suppress eccentricity excitation (Thompson, 2011; Liu et al., 2015; Anderson et al., 2017; Grishin et al., 2018a, 2020; Liu & Lai, 2021). Recently, Mangipudi et al. (2022) integrated these effects into a unified formula for the maximal eccentricity, applicable to any dynamically stable triple system with an equal-mass inner binary. In this study, we apply these results to identify orbital configurations that lead to either a stellar merger of the inner binary, or high eccentricities when the inner binary has reached a BBH configuration, resulting in a short coalescence time.
@@
The structure of this paper is as follows. First, we present a detailed stellar model of a CHE binary star system (§2.1), which we integrate with our analytical method to assess the effect of ZLK oscillations (§2.2). We proceed to analytically calculate the maximum eccentricity that a tertiary body can induce on the inner binary and evaluate how short-range forces might partially or completely suppress any increase in eccentricity (§3.1). We then determine which CHE binaries will merge as hydrogen-rich or helium-rich binaries (§3.2) and discuss the outcome of BBH that survived stellar evolution (§3.3). We discuss our results and their implications (§4) and finalize with a summary and the conclusions (§5).
@@ 2.1.1 Stellar Models
We use the open-source one-dimensional stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA; Paxton et al. 2011, 2013, 2015, 2018, 2019), version 11701, to create a model of a CHE binary. Our setup has identical physical assumptions to those of du Buisson et al. (2020), which closely aligns with the initial numerical implementation of CHE binaries presented in Marchant et al. (2016), but extends until core carbon depletion. Our default CHE binary consists of equal-mass components on a circular orbit with inner orbital period $P_{\rm{in}}=1.1$ d, each with a zero-age main-sequence (ZAMS) mass of $m_{1}=m_{2}=55\ M_{\odot}$. We choose an initial metallicity of $Z=0.00042$ (for reference, Solar metallicity is $Z_{\odot}=0.0142$, Asplund et al. 2009). We will use this CHE model to evaluate the impact of a tertiary on the massive short-period inner binary.
@@
In Fig. 1 we show the evolution of the internal structure of the stellar model, a graphic representation often referred to as a Kippenhahn diagram. In these Kippenhahn diagrams, we present the mass ($m$) and radial ($r$) coordinates as functions of time ($t$) for a component star in the CHE binary system, highlighting the evolution of the convective regions within these models. In this CHE binary, both component stars evolve identically. At the ZAMS, the $55M_{\odot}$ star has a radius of $\approx 7.6\ R_{\odot}$, and a convective core mass of $\approx 37.8M_{\odot}$ with a radius of $\approx 3.5\ R_{\odot}$. Throughout the main sequence, the star remains similarly compact, expanding up to $\approx 8.7\ R_{\odot}$. By the end of the main sequence, the total mass has decreased to $\approx 50M_{\odot}$ while the convective core has grown in mass to $\approx 46M_{\odot}$. After the main sequence, the star contracts to $\approx 2.4\ R_{\odot}$, at which point the core temperature rises sufficiently to initiate the core-helium-burning phase. Throughout this phase, and until the end of the evolution, the model slowly contracts while losing mass through stellar winds. In addition, this mass loss widens the orbit over time, leading to a final orbital period of $\approx$1.7 days by the end of the evolution. Our last model has a final mass of $\approx 43.8M_{\odot}$ and a radius of $\approx 0.9R_{\odot}$. For this orbital configuration, if we consider these systems as BBH progenitors, the time to coalescence through gravitational-wave emission would be $\approx 450$ Myr (Peters, 1964).
@@
Figure 1: Kippenhahn diagrams showing mass ($m$) and radial ($r$) coordinates as a function of time ($t$) for a 55 $M_{\odot}$ star model with metallicity $Z=0.00042$, undergoing chemically homogenous evolution. The total mass or radius is shown with a thick solid black curve. Convective regions are highlighted in purple, and regions with overshooting are marked in green. In addition, the apsidal motion constant ($k$) is shown as a red solid line, with its scale in the right side of the ordinate axis. Finally, we show a blue dashed vertical line at 5.12 Myr; we later use this model to assess stellar mergers shortly after the end of the main sequence.
@@ 2.1.2 Apsidal Motion Constant
The apsidal motion constant is crucial for evaluating the impact of external precession on the inner orbit and its competition with the ZLK mechanism (§2.2.2). Fig. 1 illustrates the apsidal motion constant ($k$) as calculated in MESA. This dimensionless number is used to parametrize the internal structure of a star and is related to the Love number ($k_{\rm{L}}$) through the relationship $2k=k_{\rm{L}}$. A lower $k$ value indicates a more centrally concentrated mass distribution of the star, which results in a slower rate of apsidal motion. For our stellar model, the value of the apsidal motion constant at the ZAMS is $k\approx 0.024$; by the end of the main sequence, this value drops to $k\approx 0.015$. After the main sequence, this value keeps decreasing; for the last saved model in our simulation $k\approx 0.0007$.
@@ 2.2 Triple Dynamics
Figure 2: Dimensionless quantities for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity $Z=0.00042$, evaluated at the zero-age main sequence. Dark grey region indicates dynamically unstable triples. We display the inner to outer orbital angular momentum ratio ($\eta:=L_{\rm{in}}/L_{\rm{out}}$, top panel) and the single-averaging parameter ($\epsilon_{\rm SA}$, bottom panel).
Figure 3: Dimensionless parameters quantifying the role of short-range forces for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity $Z=0.00042$, evaluated at the zero-age main sequence. We use different colours to highlight various regions of interest. Dark grey indicates dynamically unstable triples. We display the relative strength of the apsidal advance due to general relativity ($\epsilon_{\rm GR}$, top panel) and the relative strength of the apsidal advance due to tides ($\epsilon_{\rm Tide}$, bottom panel). For a direct comparison, we select an identical scale for the colorbars.

### v2

# Prompt stellar and binary black hole mergers in tight triples

@@
22institutetext: School of Physics and Astronomy, Monash University, Clayton, VIC 3800, Australia 33institutetext: OzGrav: Australian Research Council Centre of Excellence for Gravitational Wave Discovery, Clayton, VIC 3800, Australia 44institutetext: Institute for Astronomy, School of Physics, Zhejiang University, 310058 Hangzhou, China 55institutetext: Center for Computational Astrophysics, Flatiron Institute, New York, NY 10010, USA 66institutetext: Department of Astrophysical Sciences, Princeton University, 4 Ivy Lane, Princeton, NJ 08544, USA 77institutetext: National Astronomical Observatory of Japan, National Institutes of Natural Sciences, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan
Prompt stellar and binary black hole mergers in tight triples
Insights from chemically homogeneous evolution
A. Vigna-Gómez 11    E. Grishin 2233    J. Stegmann 11    A. Olejak 11    S. A. Popa 11    B. Liu 44    A. S. Rajamuthukumar 11    L. A. C. van Son 5566    A. Bobrick 2233    A. Dorozsmai 77
@@
Short-period massive binary stars are predicted to undergo a chemically homogeneous evolution (CHE), making them prime candidates for producing binary black holes (BBHs) that have the potential to merge within the age of the Universe. Most of these binaries have a tertiary companion and here we explore how a nearby third body could possibly influence this evolutionary channel. Our analysis combines analytic treatments of triple dynamics with insights from detailed stellar evolution models, focusing on the role of the von Zeipel-Lidov-Kozai mechanism, while also accounting for tidal and general relativistic apsidal precession. We examine the dynamics of triples at three critical evolutionary stages: the zero-age main sequence, shortly after the main sequence, and at the time of BBH formation. We find that, for triples with outer orbital periods less than 70 d(120 d), the inner binary can merge during(or after) the main sequence stage, leading to a hydrogen-rich (helium-rich) stellar merger. If a stellar merger is avoided, the inner binary could eventually form a BBH. In mildly hierarchical triples, with outer periods of around 100 d, the tertiary component can trigger a rapid merger of the BBH on timescales comparable to the outer orbital period. Stellar tides play a crucial role in determining the fate of the inner binary in such tight triple systems, as they can suppress the perturbative effects of the third star. When tidal forces damp the oscillations induced by the tertiary, the BBH merger may occur soon after stellar collapse. Notably, these outcomes are not restricted to CHE binaries but they can also be applied to any BBH formed from stars in tight orbits. Mergers in these systems are characterised by the proximity of a tertiary companion and the presence of recently ejected gas, making them promising candidates for electromagnetic counterparts and gravitational wave signals influenced by nearby tertiary objects.
@@ 1 Introduction
Stars that are several times more massive than the Sun typically form and evolve in multiple systems (e.g. Offner et al., 2023; Li et al., 2024), where they are orbited by one or more distant stellar companions. In multiple-star systems, the evolution of stars can be influenced by the interplay of mass transfer episodes, explosions, and gravitational dynamics (e.g. Eggleton, 2006; Tauris & van den Heuvel, 2023). Recent discoveries reveal that some systems initially classified as binaries are in fact part of a triple in which the most massive star is the outermost tertiary (e.g. Eisner et al., 2022; Kostov et al., 2024). In triples, the tertiary star can dynamically influence the evolution of the system. In this paper, we examine the role of a tertiary star in ‘tight’ high-mass stellar triples, where the tertiary’s outer orbital period is less than 1000 days. Our focus in this work is on binaries that experience chemically homogeneously evolution (CHE).
@@
This type of evolution has been proposed in the context of single, rapidly rotating stars, where rotationally induced mixing leads to contraction, rather than the more typical expansion observed in non-rotating stars during the main sequence (see Maeder, 1987, and reference therein). Overall, CHE is predicted to occur for (initially) massive stars ($\gtrsim 40\ M_{\odot}$) in short-orbital-period binaries ($\lesssim 3\,\rm d$) at low metallicities ($Z\ll Z_{\odot}$), where tidal forces and minimal mass loss maintain the stars rotating rapidly (e.g. de Mink et al., 2009; Song et al., 2013; Hastings et al., 2020). In addition, CHE stars are the predicted progenitors of long gamma-ray bursts (Yoon et al., 2006; Aguilera-Dena et al., 2018). CHE binaries that remain massive and compact throughout their evolution are predicted to form binary black holes (BBHs) that can merge within the age of the Universe (Marchant et al., 2016; Mandel & de Mink, 2016; de Mink & Mandel, 2016; du Buisson et al., 2020; Riley et al., 2021), making them potentially detectable through gravitational waves. CHE models predict that BBHs formed through this channel exhibit component masses $\gtrsim 17\ M_{\odot}$, nearly equal-mass ratios, and moderate–to-high black-hole spins that are aligned with the orbital angular momentum (e.g. du Buisson et al., 2020; Marchant et al., 2024). Notably, the inferred properties of BBH merger GW190517 satisfy these criteria (Qin et al., 2022) and they are consistent with a potential CHE origin. In addition, multiple stellar candidates have been proposed in the literature (see Sharpe et al., 2024, and references therein). However, there is no confirmed detection of a CHE binary from observations of stellar binaries.
@@
Given that $\sim 73\pm 16\%$ of massive ($\gtrsim 16\ M_{\odot}$) main sequence stars are found in triples and/or quadruples (Moe & Di Stefano, 2017), it is natural to consider CHE binary evolution with a tertiary companion. Tight hierarchical triples with CHE inner binaries were first studied semi-analytically in the context of gravitational-wave sources, where a coplanar triple may undergo two sequential BBH mergers that result in a peculiar, massive black-hole remnant (Vigna-Gómez et al., 2021). The CHE scenario was subsequently integrated into rapid population synthesis (Riley et al., 2021), enabling the exploration of its role within a stellar triple system (Dorozsmai et al., 2024). The rapid population synthesis of triples with a CHE binary has shed light on two particular outcomes. First, mass transfer of the tertiary onto the inner binary is common, potentially affecting the evolution of the inner binary (Dorozsmai et al., 2024; Kummer et al., 2025). Second, triple dynamics can dynamically assist a merger in the inner binary (Dorozsmai et al., 2024). A dynamically assisted merger can happen swiftly (e.g. Vigna-Gómez et al., 2022), reducing the system from a triple to a binary (e.g. Preece et al., 2024). This change impacts the number of tight inner massive binaries that would undergo CHE if they evolved in isolation.
@@
In this paper, we present an analysis of the impact of a tertiary on dynamically assisted mergers in a CHE binary. Specifically, we assess the role of the von Zeipel-Lidov-Kozai (ZLK) mechanism (von Zeipel, 1910; Lidov, 1962; Kozai, 1962) in tight triples. In the ZLK mechanism, coherent torques exchange the angular momentum between the inner and outer orbits of a hierarchical triple (Naoz, 2016). The early studies of the ZLK mechanism were done in the test-particle limit, where one member of the inner binary is nearly massless, so that the outer angular momentum significantly exceeds the inner one. In this limit, the mutual inclination required to induce a large inner eccentricity is close to 90 deg (e.g. Naoz et al., 2013; Tremaine, 2023, and references therein). However, as the ratio of the inner to the outer orbital angular momentum increases, the critical inclination varies (Anderson et al., 2017; Hamers, 2021). Systems with mild hierarchies, where the ratio of the inner and outer orbital period is not too small, can also achieve larger eccentricities (Antonini & Perets, 2012; Antognini, 2015; Luo et al., 2016; Grishin et al., 2017, 2018b). Nonetheless, additional sources of apsidal precession, such as short range forces, may suppress any eccentricity excitation (Thompson, 2011; Liu et al., 2015; Anderson et al., 2017; Grishin et al., 2018a, 2020; Liu & Lai, 2021). Recently, Mangipudi et al. (2022) integrated these effects into a unified formula for the maximal eccentricity, applicable to any dynamically stable triple system with an equal-mass inner binary. In this study, we apply these results to identify orbital configurations that lead to either a stellar merger of the inner binary, or high eccentricities when the inner binary has reached a BBH configuration, resulting in a short coalescence time.
@@
The structure of this paper is as follows. First, we present a detailed stellar model of a CHE binary star system (Sect. 2.1), which we integrate with our analytical method to assess the effect of ZLK oscillations (Sect. 2.2). We proceed to analytically calculate the maximum eccentricity that a tertiary body can induce on the inner binary and evaluate how short-range forces might partially or completely suppress any increase in eccentricity (Sect. 3.1). We then determine which CHE binaries will merge as hydrogen-rich or helium-rich binaries (Sect. 3.2) and discuss the outcome of BBH that survived stellar evolution (Sect. 3.3). We discuss our results and their implications (Sect. 4). Finally, we present a summary and our conclusions (Sect. 5).
@@ 2.1.1 Stellar models
We used the open-source one-dimensional stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA; Paxton et al. 2011, 2013, 2015, 2018, 2019), version 11701, to create a model of a CHE binary. Our setup has identical physical assumptions to those reported in du Buisson et al. (2020), which is closely aligned with the initial numerical implementation of CHE binaries presented in Marchant et al. (2016); however, the former study extends this process until core carbon depletion. Our default CHE binary consists of equal-mass components on a circular orbit with inner orbital period $P_{\rm{in}}=1.1$ d, each with a zero-age main sequence (ZAMS) mass of $m_{1}=m_{2}=55\ M_{\odot}$. The stellar models were constructed by placing two non-rotating stars in a close binary system. Once the system rapidly reached a synchronous state, we started simulating the system. We chose an initial metallicity of $Z=0.00042$ (for reference, Solar metallicity is $Z_{\odot}=0.0142$, Asplund et al. 2009). We used this CHE model to evaluate the impact of a tertiary on the massive short-period inner binary.
@@
In Fig. 1, we show the evolution of the internal structure of the stellar model, a graphic representation often referred to as a Kippenhahn diagram. In these Kippenhahn diagrams, we present the mass ($m$) and radial ($r$) coordinates as functions of time ($t$) for a component star in the CHE binary system, highlighting the evolution of the convective regions within these models. In this CHE binary, both component stars evolve identically. At the ZAMS, the $55M_{\odot}$ star has a radius of $\approx 7.6\ R_{\odot}$, and a convective core mass of $\approx 37.8M_{\odot}$ with a radius of $\approx 3.5\ R_{\odot}$. Throughout the main sequence, the star remains similarly compact, expanding up to $\approx 8.7\ R_{\odot}$. By the end of the main sequence, the total mass has decreased to $\approx 50M_{\odot}$ while the convective core has grown in mass to $\approx 46M_{\odot}$. After the main sequence, the star contracts to $\approx 2.4\ R_{\odot}$, at which point the core temperature rises sufficiently to initiate the core-helium-burning phase. Throughout this phase, and until the end of the evolution, the model slowly contracts while losing mass through stellar winds. In addition, this mass loss widens the orbit over time, leading to a final orbital period of $\approx$1.7 days by the end of the evolution. Our last model has a final mass of $\approx 43.8M_{\odot}$ and a radius of $\approx 0.9R_{\odot}$. For this orbital configuration, if we consider these systems as BBH progenitors, the time to coalescence through gravitational-wave emission would be $\approx 450$ Myr (Peters, 1964).
@@
Figure 1: Kippenhahn diagrams showing mass ($m$) and radial ($r$) coordinates as a function of time ($t$) for a 55 $M_{\odot}$ star model with metallicity $Z=0.00042$, undergoing chemically homogenous evolution. The total mass or radius is shown with a thick solid black curve. Convective regions are highlighted in purple, and regions with overshooting are marked in green. In addition, the apsidal motion constant ($k$) is shown as a solid red line, with its scale in the right side of the ordinate axis. Finally, we show a vertical dashed blue line at 5.12 Myr; we later use this model to assess stellar mergers shortly after the end of the main sequence.
@@ 2.1.2 Apsidal motion constant
The apsidal motion constant is crucial for evaluating the impact of external precession on the inner orbit and its competition with the ZLK mechanism (Sect. 2.2.2). Figure 1 illustrates the apsidal motion constant ($k$) as calculated in MESA. This dimensionless number is used to parametrise the internal structure of a star and is related to the Love number ($k_{\rm{L}}$) through the relationship $2k=k_{\rm{L}}$. A lower $k$ value indicates a more centrally concentrated mass distribution of the star, which results in a slower rate of apsidal motion. For our stellar model, the value of the apsidal motion constant at the ZAMS is $k\approx 0.024$; by the end of the main sequence, this value drops to $k\approx 0.015$. After the main sequence, this value keeps decreasing; for the last saved model in our simulation $k\approx 0.0007$.
@@ 2.2 Triple dynamics
Figure 2: Dimensionless quantities for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity $Z=0.00042$, evaluated at the ZAMS. The dark grey region indicates dynamically unstable triples. We display the inner to outer orbital angular momentum ratio ($\eta:=L_{\rm{in}}/L_{\rm{out}}$, top panel) and the single-averaging parameter ($\epsilon_{\rm SA}$, bottom panel).
Figure 3: Dimensionless parameters quantifying the role of short-range forces for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity of $Z=0.00042$, evaluated at the ZAMS. We use different colours to highlight various regions of interest. Dark grey indicates dynamically unstable triples. We display the relative strength of the apsidal advance due to GR ($\epsilon_{\rm GR}$, top panel) and the relative strength of the apsidal advance due to tides ($\epsilon_{\rm Tide}$, bottom panel). For a direct comparison, we select an identical scale for the colourbars.